derivative of x^x^x

Grade:12

1 Answers

Rinkoo Gupta
10 years ago
y=x^x^x
taking log on both sides, we get
logy=xlog(x^x)
=>logy=x.xlogx=x^2logx
differentiating w. r. to x
(1/y).dy/dx=x^2.(1/x)+logx.2x
=x+2xlogx
=x(1+2logx)
dy/dx=y.x(1+2logx)
=(x^x^x).x(1+2logx)

Thanks & Regards
Rinkoo Gupta
askIItians Faculty